OpenStudy (anonymous):

What is the distance between (7, 1) and (-4, -5)? please help

OpenStudy (anonymous):

do 7-(-4) then do 1-(-5)

OpenStudy (anonymous):

11 and 6?

OpenStudy (anonymous):

then put them in fraction

OpenStudy (anonymous):

\[3\sqrt{5}\]

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

that's not in faction mode, that's radical.. how did you get that?

OpenStudy (anonymous):

im unsure I jumped to that answer in my head and put it down

OpenStudy (anonymous):

do you have choices for the question?

OpenStudy (anonymous):

yes they are that one and \[\sqrt{137}\] \[\sqrt{157}\] 5

OpenStudy (anonymous):

do you know the distance formula

OpenStudy (anonymous):

what is here is all I know

OpenStudy (anonymous):